OpenStudy (anonymous):

Select the factors of x2 − 12x + 36. (x − 6)(x − 6) (x + 6)(x + 6) (x + 3)(x + 12) (x − 3)(x − 12)

OpenStudy (anonymous):

Ok so it has to equal 36 when multiplied but also equals -12 when added.

OpenStudy (theviper):

\[=x^2-6x-6x+36\]\[=x(x-6)-6(x-6)\]\[=(x-6)(x-6)=(x-6)^2\]

OpenStudy (anonymous):

6x6=36, 6+6=12 12x3=36, 12+3=15 -12x-3=36, -12+(-3)=-15 so its A

OpenStudy (theviper):

A. dude